3.24
The function of f(x)=8x+2 is one to one
Find an equation for f^-1(x), the inverse function.
f^-1 (x) =
(type an expression for the inverse. Use integers or fractions for any numbers in the expression.)

Answer by fractalier(6550)

To find an inverse function, follow these steps
f(x) = 8x + 2 (put in y for f(x))
y = 8x + 2 (now switch y and x)
x = 8y + 2 (now solve for y again)
x - 2 = 8y
y = (1/8)(x-2) (now put f^-1(x) in for y and you're done...)
f^-1(x) = (1/8)(x-2)